OpenStudy (anonymous):

Which is an equation in point-slope form for the given point and slope? Point: 8, -4 ; Slope: -5/6 A. y + 4 = -5/6(x + 8) B. y - 4 = -5/6(x + 8) C. y + 4 = -5/6(x - 8) D. y - 4 = -5/6(x - 8)

OpenStudy (anonymous):

b?

Nnesha (nnesha):

why not C or A ?? @Xthreefortyone

Nnesha (nnesha):

what's the formula for point slope ?? slope intercept form is y=mx+b what's the original equation for point slope form ?? :=)) do you know ?? :=))

OpenStudy (anonymous):

nope just learned this today XD

Nnesha (nnesha):

so how did you get the asnwer ?

Nnesha (nnesha):

answer*

OpenStudy (anonymous):

umm XD i guessed

OpenStudy (anonymous):

that and my 100% a Friend told me it was A

Nnesha (nnesha):

\[\huge\rm y-y_1 = m(x-x_1)\] formula for point slope form wehre m is slope and (X_1 ,y_1 ) the point where line passes through

OpenStudy (anonymous):

oh ok

Nnesha (nnesha):

so just plug in the number (x,y) in order pair first number is x-coordinate and 2nd one y now don't make guess tell me like this " i'm 100% sure my answer is ...."

OpenStudy (anonymous):

\[\huge\rm 4-4_1 = m(8-8_1)\]

OpenStudy (anonymous):

hmmm 0V1 = m(0V1) XD

Nnesha (nnesha):

hmm leave don't replace y or x

OpenStudy (anonymous):

???

Nnesha (nnesha):

\[\huge\rm y-\color{ReD}{y_1} = m(x-\color{blue}{x_1})\] \[\huge\rm (\color{blue}{x_1} ,\color{red}{y_1} ) \rightarrow (8 ,-4)\]

OpenStudy (anonymous):

ok

Nnesha (nnesha):

so just replace (x_! ,y_1) with (8,-4)

OpenStudy (anonymous):

mhm

OpenStudy (anonymous):

SO ITS A?????

OpenStudy (anonymous):

:P

Nnesha (nnesha):

hmm there is a negative sign in the formula and y value is also negative negative times negative = positive so ??

Nnesha (nnesha):

\[\huge\rm y-\color{ReD}{(-4)} = m(x-\color{blue}{(8)})\] \[\large\rm (\color{blue}{x_1} ,\color{red}{y_1} ) \rightarrow (\color{blue}{8 },\color{reD}{-4})\]

OpenStudy (anonymous):

:P so umm is it y + 4 = m(x - 8?)

OpenStudy (anonymous):

XD i just relised that

Nnesha (nnesha):

right

OpenStudy (anonymous):

because {-} (-)4 = + 4

Nnesha (nnesha):

right!

OpenStudy (anonymous):

-_- im stupid its C